-11<2x-1less than or equal to -5
Answer by haileytucki(390)

-11<2x-1<=-5
Move all terms not containing x from the center section of the interval inequality.
-11+1<2x<=-5+1
Add 1 to -11 to get -10.
-10<2x<=-5+1
Add 1 to -5 to get -4.
-10<2x<=-4
Divide each term in the inequality by 2.
-(10)/(2)<(2x)/(2)<=-(4)/(2)
Reduce the expression -(10)/(2) by removing a factor of 2 from the numerator and denominator.
-5<(2x)/(2)<=-(4)/(2)
Reduce the expression (2x)/(2) by removing a factor of 2 from the numerator and denominator.
-5 Reduce the expression -(4)/(2) by removing a factor of 2 from the numerator and denominator.
-5 Convert the solution to set notation.
(-5,-2]